OpenStudy (anonymous):

Ten times the sum of a number and 18 is at least -19.

OpenStudy (anonymous):

10(n+18)>=-19 ?

OpenStudy (gorv):

10(x+18)>=-19

OpenStudy (gorv):

x+18>=-19/10

OpenStudy (gorv):

x>=-1.9-18 x>=-19.9

OpenStudy (anonymous):

You told me the same exact thing that I wrote :s

OpenStudy (gorv):

yeahhh